# Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions

### v1

@@ 2.1 Dataset and study design
Transcriptomic features are selected to match those employed in the successful multimodal models Chen et al. [2020, 2021]. As proposed previously Chen et al. [2021], transcriptomic data for each patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors  [^2] https://github.com/mahmoodlab/MCAT/blob/master/datasets_csv_sig/signatures.csv. The number of available gene expression levels for each group are 84, 314, 498, 424, 1396 and 428 respectively for TCGA-GBMLGG resulting in a total of 3144 genes (Figure 1.a.iii). Whereas for TCGA-KIRC, the number of available gene expression levels for each group are 145, 522, 669, 681, 2366, and 990 respectively resulting in a total of 5373 genes (Figure 1.b.iii). It is to be noted that there is some overlap (by design) between the gene sets due to the intersecting functionalities they represent. Since our model processes the gene expression levels of all genes in a gene set together, the gene expression embedding corresponding to a gene group is distinct from the embeddings of the other gene groups even if the intersection of the gene sets is not null. Hence, in our paper, we comment on the importance of different gene groups to explain the grade and survival risk predictions and not on a particular gene. When discussing gene expression levels throughout this paper, we refer to transcript abundance, rather than protein levels.
Our methodology synthesizes transcriptomic data from WSIs using our diffusion based generative model, PathoGen. Synthesized transcriptomes are then used for automated cancer gradation and survival risk estimation to assess whether transcriptomic tests should be requested (Figure 1.c). We use transcriptomic data (both real and synthesized) at two stages in the predictive model (MCAT_GR) - to co-attend the WSI patch embeddings and learn a combined feature map that is suitable to predict both grade and survival and to predict the survival risk of a patient in a learnable ensemble method along with the co-attended feature maps. We perform extensive experiments using synthesized transcriptomic data to decide on this model setting (supplementary Table 1). The model is trained end-to-end for performing both gradation and survival risk estimation together using the loss function, $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$, where $L_{survival}$ is the loss for survival risk estimation, $L_{grade}$ is the loss for gradation and $\lambda$ is the loss coefficient to balance the emphasis on risk estimation and gradation. Experiments show that at $\lambda=0.3$ the model performance is optimal with close and high scores for both gradation and survival risk estimation together for both cohorts (supplementary Figure 8 and supplementary Table 1). Thus for all results reported in the paper, we use $\lambda=0.3$. For uncertainty estimation using conformal prediction, we choose the error rate $\alpha=0.1$. Further details of model architecture, loss function and training methodology are provided in Methods.
@@ 2.2 Similarity evaluation of synthesized and real transcriptomic data
The transcriptomic data synthesized by our model, PathoGen, is compared to the real patient transcriptomes to assess similarity. The Spearman correlation between real and synthesized transcriptomic features comprising all genes evaluates to 0.71 for TCGA-GBMLGG ($p-value=1.18\times 10^{-67}$) and 0.72 ($p-value=3.65\times 10^{-69}$) for TCGA-KIRC, and the mean absolute error (MAE) evaluates to 0.1 for both cohorts (Figures 2.a.i, 2.b.i and supplementary Table 2). The gene expression levels for the gene groups show no significant differences between real and synthesized transcriptomes, for both cohorts (unpaired $t$-test $p$-value $>0.05$, supplementary Table 2). Though there exists a high correlation and low MAE between synthetic and real transcriptomic data, there is some variation across the gene sets (Figures 2.a.i, 2.b.i and supplementary Table 2). Notably, synthesized and real oncogene expression levels are highly correlated across both cohorts (Figures 2.a.i, 2.b.i). This is particularly beneficial for cancer grading or survival estimation as oncogenes when mutated or overexpressed serve as biomarkers that drive tumour cell proliferation and help in predicting cancer prognosis Shay [1999]. Thus we conclude that the synthesized transcriptomic data are highly correlated with the real transcriptomic data and we gain confidence towards using the former in the predictive pipeline.
@@ 2.4.2 Visualisation of how transcriptomic data map to histopathology images
To add explainability, we compare the co-attention maps between transcriptomic data and the corresponding WSI patches, obtained using real and synthesized gene expression levels. TCGA-GBMLGG obtains a Spearman correlation of 0.939 and MAE of 2.07 and TCGA-KIRC obtains a Speraman correlation of 0.851 and MAE of 0.59 (Figures 2.a.iv, 2.b.iv and supplementary Table 3). A higher correlation with a high corresponding MAE would denote that the range of co-attention values is large for the particular gene group but the co-attention map is still correlated. For both cohorts, co-attention maps attributing to oncogenes are most correlated followed by tumour suppressor genes and transcription factor genes (Figures 2.a.iv, 2.b.iv and supplementary Table 3). Oncogenes drive cell division and survival, thus promoting cancer, inactivation of tumour suppressor genes can lead to cancer by failing to regulate uncontrolled cell growth, and transcription factor genes influence cell fate and cancer progression by regulating gene expressions, including those of tumour suppressor and oncogenes Hanahan and Weinberg [2000]. As these gene types happen to be strongly indicative of cancer diagnosis and prognosis, having co-attention maps of real and synthesized transcriptomic data to be most correlated for these groups explains the reason why the predicted grade and survival estimates using real and synthesized gene expression levels are also close.
Further, we study the contribution of synthesized transcriptomic features in co-attending to the WSI patch embeddings. We observe that there are variances in the percentage contribution of individual gene groups in co-attention across both cohorts (Figures 2.a.v,2.b.v). For both cohorts, the transcription factor genes and oncogenes have the highest co-attention values with the WSI patches for both real and synthesized gene expression levels and hence contribute the most to the prediction (Figures 3.d, 2.a.v, 4.d, 2.b.v and supplementary Figure 10.d, 11.d). Activation of oncogenes is known to drive tumour proliferation Shay [1999], hence having these as one of the higher contributors to co-attention would be aligned with clinical expectations. Also since transcription factors regulate gene expression levels to influence cancer progression, having these genes as another primary contributor in co-attention explains the reason behind our high prediction accuracy.
We zoom on specific regions of the WSI and compare the distributed grades, distributed risk and the co-attention maps between synthesized transcriptomic features and the WSI patches (Figures 3.e, 4.e and supplementary Figure 10.e, 11.e). The chosen regions are highlighted by bounding boxes of colours green, cyan and blue. We observe that true grade regions correspond to intermediate to high attention scores of transcription factors, and lower-grade regions are highly related to low attention values, as evident from the regions highlighted by the blue bounding box (Figures 3.c, 4.c) Since, transcription factors regulate the gene expression levels for tumour proliferation Hanahan and Weinberg [2000], such correlation between co-attention value and predicted grade explains the reason behind accurate predictions.
@@ 2.5.1 Quantifying uncertainty
We observe that the uncertainty profile of using synthesized transcriptomic data matches with that of using real transcriptomic data for gradation and survival risk estimation, for different patient demographics and stratification across both cohorts, by performing conformal prediction (CP) as described in Methods. For the entire test set population in TCGA-GBMLGG cohort, gradation uncertainty increases from 0.407 to 0.413 when synthesized data is used and the survival risk estimation uncertainty reduces from 0.446 to 0.438 (Figure 5.a.i). For TCGA-KIRC, gradation uncertainty increases from 0.292 to 0.334 and the survival risk uncertainty increases from 0.898 to 0.909 (Figure 6.a.i). However, the changes in uncertainty are not significant proving the similarity in uncertainty profile (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). Similarly for the different demographic groups for both cohorts, the uncertainty profile remains similar for real and synthesized transcriptomic data with no significant difference in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). It is observed that for a few population groups across both cohorts, the required coverage has not been met, because of less number of samples belonging to the calibration set. But for many such cases, using synthesized data has increased the coverage at the cost of an insignificant increase in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, Figures 6.a.i, 6.a.ii). As risk estimation directly uses the transcriptomic data for prediction and not just for co-attention, such differences in obtained coverage are expected even for minimal differences in the gene expression levels between the synthesized and real transcriptomic features. Thus, the uncertainty estimates help to understand the extent to which we may rely on the predictions, like the survival risk bounds. Additionally, the cancer grade conformal sets provide a range of alternative grades to consider before making the final decision.
@@ 2.5.2 Fairness of predictions
Next, we investigate algorithmic fairness, by assessing whether the uncertainty in gradation and survival risk predictions shows demographic biases. We observe that the predicted risk is less for patients who are reported to be alive and more for deceased patients, which is as expected (Figures 5.b, 6.b). Also, the uncertainty of survival risk estimation decreases with increasing risk, which is expected because the actual survival time of patients who are alive is beyond the reported survival time. Thus the model estimated survival risk should correspond to a time beyond the reported survival time and have a higher uncertainty measure when compared in terms of reported survival time. It may also be observed that the model tends to correlate higher risks with the highest grade, grade IV, for both cohorts. However, the trend of gradation uncertainty is different for the two cohorts. For TCGA-GBMLGG, the uncertainty of grade prediction is more for grade III, least for grade IV, and intermediate for grade II (supplementary Table 4 and Figure 5.b). But for TCGA-KIRC, the uncertainty is less for grade I and grade II with low estimated survival risks but high for all other grades (supplementary Table 4 and Figure 6.b). In general, the uncertainty of survival risk estimation is always greater if the survival time reported for a living patient is less (Figures 5.b, 6.b). Also, the uncertainty of prediction for alive patients is always high, except for gradation for TCGA-KIRC (Figures 5.a.iv, 6.a.iv). The uncertainty of grade prediction is more for male patients than females for both cohorts, whereas, for survival risk estimation, the uncertainty is more for females in TCGA-GBMLGG and males in TCGA-KIRC (Figures 5.a.iii, 6.a.iii). The uncertainty of risk estimation decreases with age for both cohorts (Figures 5.a.ii, 6.a.ii). For TCGA-GBMLGG, the uncertainty of grade prediction increases with decreasing magnification but remains similar for all magnification levels for TCGA-KIRC where magnification level 1 corresponds to the highest magnification and level 3 the lowest (supplementary Table 4). Thus these results provide an analysis from a health equity and predictive fairness point of view as to how the performance varies across patient demographics.
@@ 3.3.1 Architecture
The architecture used for gradation and risk estimation using multi-modal histopathology and synthesized transcriptomic data is MCAT Chen et al. [2021] with an additional pipeline for gradation. The modified MCAT architecture, MCAT_GR, is illustrated in Figure 7.b. The features obtained for the pathomic transformer of MCAT_GR are used for gradation by passing through a separate global attention pooling. The transcriptomic feature embeddings from the gene transformer are not used for gradation as we experiment to see that not including the transcriptomic data for grade prediction performs better. The global attention pooling performed for risk estimation and gradation are separate to decrease the dependency between the grade and risk.
@@ 4 Discussion
This work has demonstrated, through extensive experiments and state-of-the-art results on two publicly available cancer datasets from TCGA, that using generative AI to synthesize a set of transcriptomic signatures helps to improve multimodal prediction of cancer gradation and patient survival risk when combined judiciously with features learned with a vision transformer based foundation model from digital pathology images. However, in addition to the strength of innovation and the strong performance, in order to increase the trustworthiness of our PathoGen model for clinicians and pathologists, we have presented detailed supporting evidence from the point of view of transparency (mapping back the gradation and risk estimates in a distributed manner onto the WSI images for visual interpretation), reliability (uncertainty quantification by providing conformal sets for gradation and conformal bounds for risk estimation) and fairness (analysis of predictive variation across patient demographics with respect to sex, age, etc). For reproducibility of research and ethos of open science, we have made the code fully available online through our GitHub repository at https://github.com/Samiran-Dey/PathoGen.
Remarkably, we observe that the predictive performance of a multimodal model, combining PathoGen synthesized transcriptomic markers with digital pathology, is comparable to using real transcriptomic signatures. This does not replace the need for collection of transcriptomic data, since the predictive model only focuses on few pan-cancer general signatures, rather than the full transcriptome. However, this does open the door towards an exciting future line of work, that is to investigate further on whether this approach can help in estimating testing the added value of collecting transcriptomic data for the diagnostic/prognostic assessment of each patient when only routine H&\&&E stained slides are available. For future work, we envision that PathoGen can be studied as a screening tool, to guide the targeted collection of expensive transcriptomic data modalities, by potentially selecting patients with highest correlation with the predictive output of the model. This would of course need a dedicated follow up study to validate which would include patient response to treatment when selected in this manner.

### v2

@@ 2.1 Dataset and study design
Transcriptomic features are selected to match those employed in the successful multimodal models Chen et al. [2020, 2021]. As proposed previously Chen et al. [2021], transcriptomic data for each patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors  [^2] https://github.com/mahmoodlab/MCAT/blob/master/datasets_csv_sig/signatures.csv. The number of available gene expression levels for each group are 84, 314, 498, 424, 1396 and 428 respectively for TCGA-GBMLGG resulting in a total of 3144 genes (Figure 1.a.iii). Whereas for TCGA-KIRC, the number of available gene expression levels for each group are 145, 522, 669, 681, 2366, and 990 respectively resulting in a total of 5373 genes (Figure 1.b.iii). When discussing gene expression levels throughout this paper, we refer to transcript abundance, rather than protein levels.
Our methodology synthesizes transcriptomic data from WSIs using our diffusion based generative model, PathoGen. Synthesized transcriptomes are then used for automated cancer gradation and survival risk estimation to assess whether transcriptomic tests should be requested (Figure 1.c). We use transcriptomic data (both real and synthesized) at two stages in the predictive model (MCAT_GR) - to co-attend the WSI patch embeddings and learn a combined feature map that is suitable to predict both grade and survival and to predict the survival risk of a patient in a learnable ensemble method along with the co-attended feature maps. We perform extensive experiments using synthesized transcriptomic data to decide on this model setting (supplementary Table 1). The model is trained end-to-end for performing both gradation and survival risk estimation together using the loss function, $L=\lambda\times L_{survival}+(1-\lambda)\times L_{grade}$, where $L_{survival}$ is the loss for survival risk estimation, $L_{grade}$ is the loss for gradation and $\lambda$ is the loss coefficient to balance the emphasis on survival risk estimation and gradation. Experiments show that at $\lambda=0.3$ the model performance is optimal with close and high scores for both gradation and survival risk estimation together for both cohorts (supplementary Figure 8 and supplementary Table 1). Thus for all results reported in the paper, we use $\lambda=0.3$. For uncertainty estimation using conformal prediction, we choose the error rate $\alpha=0.1$. Further details of model architecture, loss function and training methodology are provided in Methods.
@@ 2.2 Similarity evaluation of synthesized and real transcriptomic data
The transcriptomic data synthesized by our model, PathoGen, is compared to the real patient transcriptomes to assess similarity. The Spearman correlation between real and synthesized transcriptomic features comprising all genes evaluates to 0.71 for TCGA-GBMLGG ($p-value=1.18\times 10^{-67}$) and 0.72 ($p-value=3.65\times 10^{-69}$) for TCGA-KIRC, and the mean absolute error (MAE) evaluates to 0.1 for both cohorts (Figures 2.a.i, 2.b.i and supplementary Table 2). The gene expression levels for the gene groups show no significant differences between real and synthesized transcriptomes, for both cohorts (unpaired $t$-test $p$-value $>0.05$, supplementary Table 2). Though there exists a high correlation and low MAE between synthetic and real transcriptomic data, there is some variation across the gene sets (Figures 2.a.i, 2.b.i and supplementary Table 2). Notably, synthesized and real oncogene expression levels are highly correlated across both cohorts (Figures 2.a.i, 2.b.i). This is particularly beneficial for cancer grading or survival estimation as oncogenes when aberrantly active can serve as biomarkers that drive tumour progression, facilitating the prediction of cancer prognosis Shay [1999]. Thus we conclude that the synthesized transcriptomic data are highly correlated with the real transcriptomic data and we gain confidence towards using the former in the predictive pipeline.
@@ 2.4.2 Visualisation of how transcriptomic data map to histopathology images
To add explainability, we compare the co-attention maps between transcriptomic data and the corresponding WSI patches, obtained using real and synthesized gene expression levels. TCGA-GBMLGG obtains a Spearman correlation of 0.939 and MAE of 2.07 and TCGA-KIRC obtains a Speraman correlation of 0.851 and MAE of 0.59 (Figures 2.a.iv, 2.b.iv and supplementary Table 3). A higher correlation with a high corresponding MAE would denote that the range of co-attention values is large for the particular gene group but the co-attention map is still correlated. For both cohorts, co-attention maps attributing to oncogenes are most correlated followed by tumour suppressor genes and transcription factor genes (Figures 2.a.iv, 2.b.iv and supplementary Table 3). Oncogenes usually drive cell division and survival, thus promoting cancer, inactivation of tumour suppressor genes can lead to cancer by failing to regulate uncontrolled cell growth Hanahan and Weinberg [2000], so the general co-attention observations above are not unexpected, although a rigorous inference would need a detailed analysis of gene set functional significance, which is outside the scope of this work.
Further, we study the contribution of synthesized transcriptomic features in co-attending to the WSI patch embeddings. We observe that there are variances in the percentage contribution of individual gene groups in co-attention across both cohorts (Figures 2.a.v,2.b.v). For both cohorts, the transcription factor genes and oncogenes have the highest co-attention values with the WSI patches for both real and synthesized gene expression levels and hence contribute the most to the prediction (Figures 3.d, 2.a.v, 4.d, 2.b.v and supplementary Figure 10.d, 11.d). Activation of oncogenes is known to drive tumour proliferation Shay [1999], hence having these as one of the higher contributors to co-attention would be aligned with clinical expectations.
We next visualise specific regions of the WSI and compare the distributed grades, distributed risk and the co-attention maps between synthesized transcriptomic features and the WSI patches (Figures 3.e, 4.e and supplementary Figure 10.e, 11.e). We observe that ground truth grade regions correspond to intermediate to high attention scores of transcription factors, and lower-grade regions are highly related to low attention values, as evident from the regions highlighted by the blue bounding box (Figures 3.c, 4.c).
@@ 2.5.1 Quantifying uncertainty
We observe that the uncertainty profile of using synthesized transcriptomic data matches with that of using real transcriptomic data for gradation and survival risk estimation, for different patient demographics and stratification across both cohorts, by performing conformal prediction (CP) as described in Methods. For the entire test set population in TCGA-GBMLGG cohort, gradation uncertainty increases from 0.407 to 0.413 when synthesized data is used and the survival risk estimation uncertainty reduces from 0.446 to 0.438 (Figure 5.a.i). For TCGA-KIRC, gradation uncertainty increases from 0.292 to 0.334 and the survival risk uncertainty increases from 0.898 to 0.909 (Figure 6.a.i). However, the changes in uncertainty are not significant proving the similarity in uncertainty profile (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). Similarly for the different demographic groups for both cohorts, the uncertainty profile remains similar for real and synthesized transcriptomic data with no significant difference in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). It is observed that for a few population groups across both cohorts, the required coverage has not been met, because of fewer samples in the calibration set. But for many such cases, using synthesized data has increased the coverage at the cost of an insignificant increase in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, Figures 6.a.i, 6.a.ii). As risk estimation directly uses the transcriptomic data for prediction and not just for co-attention, such differences in obtained coverage are expected even for minimal differences in the gene expression levels between the synthesized and real transcriptomic features. Thus, the uncertainty estimates help to understand the extent to which we may rely on the predictions, like the survival risk bounds. Additionally, the cancer grade conformal sets provide a range of alternative grades to consider before making the final decision.
@@ 2.5.2 Fairness of predictions
Next, we investigate algorithmic fairness, by assessing whether the uncertainty (normalised prediction set size) in gradation and survival risk predictions shows demographic biases. We observe that the predicted risk is less for patients who are reported to be alive and more for deceased patients, which is as expected (Figures 5.b, 6.b). Also, the uncertainty of survival risk estimation decreases with increasing risk, which is expected because the actual survival time of patients who are alive is beyond the reported survival time. Thus the model estimated survival risk should correspond to a time beyond the reported survival time and have a higher uncertainty measure when compared in terms of reported survival time. It may also be observed that the model tends to correlate higher risks with the highest grade, grade IV, for both cohorts. However, the trend of gradation uncertainty is different for the two cohorts. For TCGA-GBMLGG, the uncertainty of grade prediction is more for grade III, least for grade IV, and intermediate for grade II (supplementary Table 4 and Figure 5.b). But for TCGA-KIRC, the uncertainty is less for grade I and grade II with low estimated survival risks but high for all other grades (supplementary Table 4 and Figure 6.b). In general, the uncertainty of survival risk estimation is always greater if the survival time reported for a living patient is less (Figures 5.b, 6.b). Also, the uncertainty of prediction for alive patients is always high, except for gradation for TCGA-KIRC (Figures 5.a.iv, 6.a.iv). The uncertainty of grade prediction is more for male patients than females for both cohorts, whereas, for survival risk estimation, the uncertainty is more for females in TCGA-GBMLGG and males in TCGA-KIRC (Figures 5.a.iii, 6.a.iii). The uncertainty of risk estimation decreases with age for both cohorts (Figures 5.a.ii, 6.a.ii). For TCGA-GBMLGG, the uncertainty of grade prediction increases with decreasing magnification but remains similar for all magnification levels for TCGA-KIRC where magnification level 1 corresponds to the highest magnification and level 3 the lowest (supplementary Table 4). Thus these results provide an analysis from a health equity and predictive fairness point of view as to how the performance varies across patient demographics.
@@ 3.3.1 Architecture
The architecture used for gradation and risk estimation using multi-modal histopathology and synthesized transcriptomic data is MCAT Chen et al. [2021] with an additional pipeline for gradation. The modified MCAT architecture, MCAT_GR, is illustrated in Figure 7.b. The co-attended features processed using the pathomic transformer of MCAT_GR are used for gradation by passing through a separate global attention pooling. However, transcriptomic feature embeddings from the gene transformer are not used directly for gradation as we experiment to see that not including the transcriptomic data for grade prediction performs better. The global attention pooling performed for risk estimation and gradation are separate to decrease the dependency between the grade and risk.
@@ 4 Discussion
Here we present PathoGen, a crossmodal generative diffusion model for synthesizing transcriptomic features from WSIs of H&\&&E stained slides. We demonstrated that PathoGen synthesized transcriptomic features are highly correlated to true patient transcriptomes. By using a multimodal AI model to predict cancer grade and survival risk from WSIs and transcriptomic features, with state-of-the-art performance, we demonstrated that combining the features learnt from WSIs using a vision transformer based foundation model with the synthesized transcriptomic features from PathoGen, significantly improved predictive accuracy compared to WSIs alone. Remarkably, we found that there was no significant difference in accuracy when real patient transcriptomic data were used instead of PathoGen synthesized transcriptomic data in the multimodal model, demonstrating that PathoGen captures the independent gradation/prognostic information in true patient gene expressions.
H&\&&E stained slides of tumour tissue are essential for diagnostic and prognostic assessment of malignancy, and are collected routinely. However, transcriptomic tests, though they have the potential to inform prognosis and guide treatment selection Sparano et al. [2018], Tsakiroglou et al. [2023] are not routinely performed, especially in the public sector Harnan et al. [2019]. This is due to a number of reasons, including economic cost, infrastructural requirements and incomplete knowledge about patient benefit. PathoGen can synthesize realistic and informative patient transcriptomic features, from routine and inexpensive WSIs of H&\&&E stained slides. Thus in the future, we envision that PathoGen can be studied as a low cost screening tool, to guide the targeted collection of transcriptomic data modalities, by potentially selecting patients who would benefit from transcriptomic assessment for more accurate diagnosis/prognosis. This would of course need a dedicated follow up study which would include patient response to treatment when selected in this manner, that would entail a separate clinical validation project.
For PathoGen to be useful in clinic, it must be trusted. To facilitate this trust we have undertaken two measures. Firstly, by employing co-attention tools for model transparency, we map gradation and survival risk estimates back onto the WSI images in a distributed manner. This has provided a window into intra-tumour heterogeneity, and can highlight cases where whole slide assessment and distributed assessment of tumour grade may differ. Secondly, we have employed conformal prediction for assessment of model reliability, providing not simply point estimates for gradation and survival risk, but ranges which contain true values with a specified certainty. This analysis has also allowed us to assess model fairness, by comparing model reliability across patient demographics such as age and gender. Importantly, we found that there was no significant difference in model reliability or fairness when using PathoGen synthesized transcriptomic features compared to real patient transcriptomes. For reproducibility of research and ethos of open science, we have made the code fully available online through our GitHub repository at https://github.com/Samiran-Dey/PathoGen.
Our initial results are based on assessment of two publicly available cancer datasets from TCGA, describing glial and renal tumours. Though it is encouraging that PathoGen shows state-of-the-art performance across diverse tumour types, further assessment in other malignancies and in independent datasets, would expand the remit of our results. PathoGen is able to reliably synthesize transcriptomic embeddings, describing 6 distinct gene sets. Though these gene sets (both synthesized and real) contain important diagnostic/prognostic information, they do not capture the full diversity of the transcriptome. There is likely a limit to how much of the full transcriptome can be synthesized from only knowledge of WSIs, and hence we do not claim that PathoGen replaces the need for real transcriptomic data, but rather it provides a potential way of estimating the benefit of collecting the full data across the population. The increasing use of techniques such as spatial transcriptomics Williams et al. [2022] will provide insight into this question, as well as WSI patch level resolution of transcriptomics, to guide development of synthesized transcriptomic data with a greater number of features in the near future.